# SIGMA: An Efficient Heterophilous Graph Neural Network with Fast Global Aggregation

### v4

@@ VI-D Components Evaluation
TABLE VIII: Component study on the effect of SIGMA and GloGNN components $\mathbf{S}$, $\mathbf{S\cdot A}$, $\mathbf{A}$, and $\mathbf{X}$. Avg.$\downarrow$ and Max.$\downarrow$ denote the average and maximum accuracy drop of the model respectively.
Component genius arxiv penn94 twitch snap pokec avg.$\downarrow$ max.$\downarrow$ SIGMA 91.68 55.16 86.31 67.21 64.63 82.33 - - SIGMA w/o $\mathbf{S}$ 87.21 53.67 83.72 63.81 61.96 81.78 2.51 3.73 SIGMA w/ $\mathbf{S\cdot A}$ 90.37 53.89 84.59 64.56 62.73 81.27 1.65 2.65 SIGMA w/o $\mathbf{X}$ 73.84 53.10 82.46 62.83 62.34 81.92 4.78 17.2 SIGMA w/o $\mathbf{A}$ 86.08 34.97 75.36 60.01 36.26 60.16 15.4 25.7 GloGNN 0.9066 0.5468 0.8557 0.6619 0.6209 0.9300 - - GloGNN w/o $\mathbf{A}$ 0.7463 0.3705 0.7978 0.5987 0.3066 0.6220 16.3 31.43 GloGNN w/o $\mathbf{X}$ 0.8718 0.5445 0.8203 0.6559 0.6153 0.8203 1.5 3.53
@@
    Global aggregation functions are critical. The accuracy decline observed with SIGMA w/ $\mathbf{S}\cdot\mathbf{A}$ emphasizes the importance of global aggregation capabilities, such as those facilitated by SimRank. This aligns with our theoretical interpretation in Section IV.3, which identifies the importance of long-distance node pairs enriched with additional homophily information. Restricting nodes detected through $S$ to immediate neighbors significantly reduces performance.
@@
VI-E Grouping Effect Visualization
To demonstrate the capabilities of our model, we present the output embedding matrix Z (Eq. (6)) and show its application on six small-scale graphs in Fig.8. As shown in Theorem IV.4, the Z matrix exhibits the desired grouping effect, confirmed visually in Fig.8. Nodes with the same category labels have similar embedding patterns, while nodes from different categories show clear distinctions. For example, in the Pubmed dataset (Fig.8), which has three node types, the embeddings form distinct rectangular patterns. The number of patterns corresponds to the number of categories, demonstrating SIGMA’s ability to capture global homophily and distinguish heterophily. These observations provide strong evidence of SIGMA’s effectiveness in capturing complex graph structures.
@@ VI-E Grouping Effect Visualization
VI-F Iterative Aggregation Mechanism Exploration
In the previous sections, SIGMA used a one-time aggregation mechanism with the SimRank matrix S, which, as shown in experiments, provides excellent performance. However, SIGMA can also be adapted to iterative computations. In Eq. (5), we perform a one-time MLP transformation and aggregation, but this design can be extended to other architectures and propagation schemes as a general edge rewiring method. For example, in an iterative GCN model, the aggregation steps are represented as: $\textbf{Z}=\sigma(...\sigma(\textbf{A}\ \sigma(\textbf{AXW})\textbf{W})...)$. By replacing A with S, the aggregation process becomes: $\textbf{Z}=\sigma(...\sigma(\textbf{S}\ \sigma(\textbf{SXS})\ \textbf{W})...)$, where $\mathbf{\mathbf{}X_{S}}=\delta\mathbf{AW_{A}}+(1-\delta)\mathbf{XW_{X}}$. We conducted exploratory experiments on SIGMA with iterative designs, as shown in Table XI. The results show that compared to GCN, SIGMA aggregation significantly improves performance, especially on heterophilous graphs, due to the combination of high-quality initial features $\mathbf{X_{S}}$ and the global aggregation from $S$. Omitting either component results in lower performance. Therefore, we consider SIGMA a unique GNN model efficiently addressing heterophily graph problems.
VII Conclusion

### v5

@@ V-D Components Evaluation
[…]
@@
    Global aggregation functions are critical. The accuracy decline observed with SIGMA w/ $\mathbf{S}\cdot\mathbf{A}$ emphasizes the importance of global aggregation capabilities, such as those facilitated by SimRank. This aligns with our theoretical interpretation in Section III.3, which identifies the importance of long-distance node pairs enriched with additional homophily information. Restricting nodes detected through $S$ to immediate neighbors significantly reduces performance.
@@
V-E Grouping Effect Visualization
To demonstrate the capabilities of our model, we present the output embedding matrix Z (Eq. (6)) and show its application on six small-scale graphs in Fig.8. As shown in Theorem III.4, the Z matrix exhibits the desired grouping effect, confirmed visually in Fig.8. Nodes with the same category labels have similar embedding patterns, while nodes from different categories show clear distinctions. For example, in the Pubmed dataset (Fig.8), which has three node types, the embeddings form distinct rectangular patterns. The number of patterns corresponds to the number of categories, demonstrating SIGMA’s ability to capture global homophily and distinguish heterophily. These observations provide strong evidence of SIGMA’s effectiveness in capturing complex graph structures.
@@ V-E Grouping Effect Visualization
V-F Iterative Aggregation Mechanism Exploration
[…]
VI Conclusion